Calculate Ksp of a salt, M2S, where M is a +1 metal. It's solubility is 5.13 x 10-6 M
Answer #1

Mas ant + 82 ass Ksp = (maj? [82] = (25) *$ = 45° Solubility = S = 5013* 16569. Ksp = UX (5.13x106) 13 540x1018 - ksp = 5.408

Thank you ?

All the best....